A train travelled 720km in 9hours. Find
(A)how far would travel in 11hours
(B)how long it would take to go 1120km

Answer by Alan3354(69443)

A train travelled 720km in 9hours. Find
(A)how far would travel in 11hours
----
Find the speed.
To find km/hr, divide the kms by the hours.
Is that not obvious?
---
d = r*t
t = 11 hours.
========================
(B)how long it would take to go 1120km
d = r*t
t = d/r